OpenStudy (anonymous):

I Need Help With Number 19!

OpenStudy (softballgirl372015):

love the elmo :)

OpenStudy (anonymous):

Question?

OpenStudy (softballgirl372015):

he is probably writing it now...

OpenStudy (softballgirl372015):

i think the answer is b

OpenStudy (anonymous):

19. DE II AB. Hence <DCE = <ACB, which is pretty much the same angle, and <CDE = <CAB, corresponding angles are equal. Hence Tri. DCE is similar to ACB. by AA. So corresponding sides are proportional; if u can hence find out X and then BC

OpenStudy (softballgirl372015):

you have to set up proportions since the two triangles are similar.

OpenStudy (anonymous):

and there is a VERY simple way to write the proportional sides. suppose 123 similar to 456, u can write 12/45 = 23/56 = 13/46 (check the corresponding characters.) same way u can write it here and evaluate. cheers! :)

OpenStudy (softballgirl372015):

\[\frac{ x+3 }{ 15}=\frac{ 2x+2 }{ 24 }\]

OpenStudy (softballgirl372015):

you then have to solve for x and then plug x back into (x+3) and (x-1) and add the two nnumbers and then you get 16

OpenStudy (softballgirl372015):

does this help?